Bill O’Brien would like to take his wife, Mary, on a trip three years from now to Europe to celebrate their 40th anniversary. He has just received a $26,000 inheritance from an uncle and intends to invest it for the trip. Bill estimates the trip will cost $33,280. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

What interest rate, compounded annually, must Bill earn to accumulate enough to pay for the trip? (Round your interest rate to the nearest whole percentage.)
Answer #1

Interest rate will be 9%

Working;

A = P*(1+i)^n Where,
$ 33,280.00 = 26000*(1+i)^3 A Future value = $ 33,280.00
1.28 = (1+i)^3 P Current value = $ 26,000.00
1.28 ^(1/3) = 1+i i Interest rate = ?
1.08576705 = 1+i n Time = 3
0.08576705 = i
So, Interest rate will be 9%